# The radius of statistical efficiency

### v1

@@ Appendix C Proofs from Section 3
We first argue the inclusion $\supset$. Observe that for any measure $\mu\in\mathcal{P}_{2}^{\circ}$ with $\lambda_{1}(\Sigma_{\mu})=\lambda_{2}(\Sigma_{\mu})$, the set of maximizers of (16) is the intersection of a sphere and the top eigenspace of $\Sigma_{\mu}$. Since none of the maximizers are isolated, they are unstable and therefore $\mu$ lies in $\mathcal{E}$.
To see the reverse inclusion $\subset$, fix a measure $\mu\in\mathcal{P}_{2}^{\circ}$ and suppose that the top two eigenvalues $\lambda_{1}$ and $\lambda_{2}$ of $\Sigma_{\mu}$ are distinct. Then the normalized top eigenvector $v$ of $\Sigma_{\mu}$ is the unique maximizer of (16). It remains to verify that $v$ is a tilt-stable maximizer of (16). To this end, define the Lagrangian function $\mathcal{L}(v,\lambda)=-\frac{1}{2}\operatorname{\mathbb{E}}_{\mu}\langle x,v\rangle^{2}+\frac{\lambda}{2}(\|v\|^{2}-1)$. Then equalities hold:
@@
In particular, the first equation shows that the optimal Lagrange multiplier $\lambda$ is $\lambda_{1}$. Let $V\in\mathbf{R}^{d\times(d-1)}$ be a matrix with an eigenbasis for $v^{\perp}$ as its columns. Then an elementary computation yields
$\min_{y\in\mathbb{S}^{d-2}}\langle\nabla^{2}_{vv}\mathcal{L}(v,\lambda)(Vy),(Vy)\rangle=\min_{y\in\mathbb{S}^{d-2}}\sum_{i=2}^{d}(\lambda_{1}-\lambda_{i})y_{i}^{2}=\lambda_{1}-\lambda_{2}>0,$ (90)
and therefore $v$ is a tilt-stable maximizer of (16). Moreover, the claimed equality ${\rm REG}(\mu)^{-1}=\lambda_{1}(\Sigma_{\mu})-\lambda_{2}(\Sigma_{\mu})$ follows directly from (90), thereby completing the proof.
@@ C.2 Proof of Theorem 3
This follows directly by applying Theorem A.2 with $\mathcal{G}=\mathcal{E}$ from Lemma 3.1 and $G=\{v\in\mathbf{R}^{d}_{+}:v^{(1)}=v^{(2)}\}$, where $v^{(i)}$ denotes the $i$’th largest coordinate value of $v$.
@@
Henceforth, fix a measure $\mu\in\mathcal{P}_{2}^{\circ}$ and define the shorthand $\lambda:=\lambda(\Sigma_{\mu})$. Let’s dispense first with the simple direction $\supset$. To this end, suppose that equality $\lambda_{q}=\lambda_{q+1}$ holds. Then we may form two sets of orthonormal bases $U:=\{u_{1},\ldots,u_{q}\}$ and $U^{\prime}=\{u_{1},\ldots,u_{q-1},u_{q}^{\prime}\}$ with $\langle u_{q},u_{q}^{\prime}\rangle=0$, and which are contained in the span of the eigenspaces corresponding to the top $q$ eigenvalues. We may further interpolate between the two bases with $U_{t}=\{u_{1},\ldots,tu_{q}+(1-t)u_{q}^{\prime}\}$ for $t\in(0,1)$. The orthogonal projections onto the span of $U_{t}$ furnish a path of optimal solutions, which are therefore not tilt-stable. Thus $\mu$ lies in $\mathcal{E}$, as claimed.
@@
We now establish the reverse inclusion $\subset$. Suppose therefore that $\lambda_{q}$ and $\lambda_{q+1}$ are distinct. We begin by conveniently parameterizing the Grassmannian manifold ${\rm Gr}(q,d)$ as follows. Define the matrix $A:=\begin{bmatrix}I_{q}&0\\ 0&0\end{bmatrix}$. Then using [6, Section 2.1] we may write ${\rm Gr}(q,d)$ as the orbit of $A$ under conjugation by orthogonal matrices:
${\rm Gr}(q,d)=\{UAU^{\top}:U\in O(d)\}.$
@@
Fix a skew symmetric matrix $W:=\begin{bmatrix}W_{1}&W_{2}\\ -W_{2}^{\top}&W_{4}\end{bmatrix}$ and define the curve $\gamma\colon\mathbf{R}\to{\rm Gr}(q,d)$ by
$\gamma(t)=\exp(-tW)\cdot A\cdot\exp(tW).$
Differentiating the curve yields the expression
$\dot{\gamma}(0)=AW-WA=\begin{bmatrix}0&W_{2}\\ W_{2}^{\top}&0\end{bmatrix}.$
@@
Moreover, [6, Section 2.3] shows that varying $W$ among all skew-symmetric matrices yields the entire tangent space
$T_{{\rm Gr}(q,d)}(A)=\left\{\begin{bmatrix}0&B\\ B^{\top}&0\end{bmatrix}:B\in\mathbf{R}^{q\times(d-q)}\right\}.$
@@
Now without loss of generality, we may assume that $\Sigma_{\mu}$ is diagonal, that is $\Sigma_{\mu}={\rm Diag}(\lambda)$. Then clearly, $R=A$ is the unique maximizer of the problem (19). We now perform the second order expansion
$(f\circ\gamma)(t)=\langle\gamma(t),\Sigma_{\mu}\rangle=f(A)+t\underbrace{\langle AW-WA,\Sigma_{\mu}\rangle}_{=0}+t^{2}\left\langle\frac{1}{2}(AW^{2}+W^{2}A)-WAW,\Sigma_{\mu}\right\rangle+O(t^{3}).$
In particular, we deduce
$(f\circ\gamma)^{\prime\prime}(0)=\left\langle\frac{1}{2}(AW^{2}+W^{2}A)-WAW,\Sigma_{\mu}\right\rangle.$
Taking into account the definition of $A$, a quick computation show
$\frac{1}{2}(AW^{2}+W^{2}A)-WAW$ $=\begin{bmatrix}W_{1}^{2}-W_{2}W_{2}^{\top}&\frac{1}{2}(W_{2}W_{4}+W_{1}W_{2})\\ -\frac{1}{2}(W_{2}W_{4}+W_{1}W_{2})^{\top}&0\end{bmatrix}-\begin{bmatrix}W_{1}^{2}&W_{1}W_{2}\\ -W_{2}^{\top}W_{1}&-W_{2}^{\top}W_{2}\end{bmatrix}$
$=\begin{bmatrix}-W_{2}W_{2}^{\top}&\frac{1}{2}(W_{2}W_{4}-W_{1}W_{2})\\ -\frac{1}{2}(W_{2}W_{4}-W_{1}W_{2})^{\top}&W_{2}^{\top}W_{2}\end{bmatrix}.$
@@
Taking the trace product with the diagonal matrix $\Sigma_{\mu}$ yields
$(f\circ\gamma)^{\prime\prime}(0)$ $=-\langle W_{2}W_{2}^{\top},{\rm Diag}(\lambda_{1:q})\rangle+\langle W_{2}^{\top}W_{2},{\rm Diag}(\lambda_{q+1:d})\rangle$ (91)
$\leq-\langle W_{2}W_{2}^{\top},\lambda_{q}I_{q}\rangle+\langle W_{2}^{\top}W_{2},\lambda_{q+1}I_{d-q}\rangle$ (92)
$=-(\lambda_{q}-\lambda_{q+1})\cdot\|W_{2}\|^{2}_{F}$ (93)
@@
In particular, the covariant Hessian $\nabla^{2}_{\mathcal{M}}f(A)$ is negative definite on the tangent space $T_{{\rm Gr}(q,d)}(A)$. Therefore, $A$ is a tilt-stable maximizer of the problem, as we had to show. Moreover, setting $W_{2}=\operatorname{Diag}(w)$ with $w_{q}=w_{q+1}=1$ and $w_{i}=0$ for $i\notin\{q,q+1\}$, yields equality in (92). Therefore we deduce ${\rm REG}(\mu)^{-1}=\lambda_{q}(\Sigma_{\mu})-\lambda_{q+1}(\Sigma_{\mu})$ as claimed.
@@ C.4 Proof of Theorem 3
This follows directly by applying Theorem A.2 with $\mathcal{G}=\mathcal{E}$ from Lemma 3.3 and $G=\{v\in\mathbf{R}^{d}_{+}:v^{(q)}=v^{(q+1)}\}$, where $v^{(i)}$ denotes the $i$’th largest coordinate value of $v$.
@@ Appendix D Proofs from Section 4
For any $\mu\in\mathcal{Q}$, observe
$\nabla f(\beta)=\operatorname{\mathbb{E}}\left[\left(h^{\prime}(\langle x,\beta\rangle)-y\right)x\right]$ $=\operatorname{\mathbb{E}}\left[\operatorname{\mathbb{E}}\left[h^{\prime}(\langle x,\beta\rangle)-y\mid x\right]x\right]$
$=\operatorname{\mathbb{E}}\left[(h^{\prime}(\langle x,\beta\rangle)-h^{\prime}(\langle x,{\beta}_{\star}\rangle))x\right]$
Therefore, equality $\nabla f({\beta}_{\star})=0$ holds for any distribution of $x$. Hence, ${\beta}_{\star}$ is critical for the problem with zero Lagrange multipliers $\lambda=0$. Differentiating again yields the expression for the Hessian
$H:=\nabla^{2}f({\beta}_{\star})=\sigma^{-2}\cdot\operatorname{\mathbb{E}}\big{[}h^{\prime\prime}(\langle x,\beta_{\star}\rangle)xx^{\top}\big{]}.$ (94)
Note that $H$ is positive semidefinite since $h^{\prime\prime}>0$. Consequently, the set of ill-conditioned distributions $\mathcal{E}$ corresponds to those distributions on $x$ for which $\ker(H)$ nontrivially intersects $\mathcal{T}$. Clearly, a vector $v$ lies in $\ker(H)$ if and only if $0=\langle Hv,v\rangle$, or equivalently $0={\operatorname{\mathbb{E}}}\big{[}h^{\prime\prime}(\langle x,{\beta}_{\star}\rangle)\langle x,v\rangle^{2}\big{]}$. Taking into account the assumption $h^{\prime\prime}>0$, this occurs precisely when $v$ lies in the nullspace of $\Sigma_{\mu}$. Thus $\mathcal{E}$ consists of all measures $\mu\in\mathcal{Q}$ satisfying $\mathcal{T}\cap\ker\left(\Sigma_{\mu}\right)\neq\{0\}.$ Finally, it follows directly from (94) that if for some $\alpha,\beta>0$ the inequality $\alpha\leq h^{\prime\prime}(\langle x,\beta_{\star}\rangle)\leq\beta$ holds for $\mu$-almost every $x$, then $\lambda_{\min}(\Sigma_{\mu}\mid_{{\mathcal{T}}})\in\lambda_{\min}\left(\Sigma_{\mu}|_{{\mathcal{T}}}\right)\cdot[\alpha,\beta]$, thereby completing the proof.
@@ D.2 Proof of Theorem 4
We begin by verifying the claim for the formulation (26). To this end, a quick computation shows $\nabla f(\beta_{\star})=0$ and $\nabla^{2}f(\beta_{\star})={\operatorname{\mathbb{E}}}_{\mu}\Big{[}\langle x,{\beta}_{\star}\rangle^{2}xx^{\top}\Big{]}.$ Therefore we deduce
@@
In particular, $\nabla^{2}f(\beta_{\star})$ if and only if the support of $\mu$ is contained in ${\beta}_{\star}^{\perp}\cup v^{\perp}$ almost surely.
Next, we verify the claim for the formulation (25). To this end, let $\mathcal{M}$ denote the set of symmetric PSD rank one matrices:
=$\mathcal{M}=\{M\in\bm{S}^{d}_{+}:\operatorname{rank}(M)=1\}.$
A quick computation now yields
$\nabla f(M)=\operatorname{\mathbb{E}}\langle M-M_{\star},xx^{\top}\rangle\qquad\textrm{and}\qquad\nabla^{2}f(M_{\star})[\Delta,\Delta]=\operatorname{\mathbb{E}}\langle\Delta,xx^{\top}\rangle^{2}.$
In particular, equality $\nabla f(M_{\star})=0$ holds and therefore the optimal Lagrange multipliers $\lambda_{\star}$ are zero. Hence the Hessian of the Lagrangian $\nabla^{2}\mathcal{L}(M_{\star},\lambda_{\star})$ coincides with $\nabla^{2}f(M_{\star})$. Classically, if we form the factorization $M_{\star}={\beta}_{\star}{\beta}_{\star}^{\top}$, then the tangent space to $\mathcal{M}$ at $M_{\star}$ can be written as
$\mathcal{T}=\{{\beta}_{\star}v^{\top}+v{\beta}_{\star}^{\top}:v\in\mathbf{R}^{d}\}.$
@@
$\nabla^{2}f(M_{\star})[\Delta,\Delta]=\operatorname{\mathbb{E}}\langle\Delta,xx^{\top}\rangle^{2}=4\operatorname{\mathbb{E}}\langle{\beta}_{\star},x\rangle^{2}\langle v,x\rangle^{2}.$
Note that $\|\Delta\|_{F}^{2}=2({\beta}_{\star}^{\top}v)^{2}+2\|{\beta}_{\star}\|^{2}\|v\|^{2}$ and therefore $\|\Delta\|_{F}^{2}/\|{\beta}_{\star}\|^{2}\|v\|^{2}\in[2,4]$. Therefore $\Delta$ is nonzero as long as $w$ is nonzero. We therefore again deduce that $\nabla^{2}f(M_{\star})[\Delta,\Delta]=0$ of some nonzero $\Delta\in T_{\mathcal{M}}(M_{\star})$ if and only if the support of $\mu$ is contained in ${\beta}_{\star}^{\perp}\cup v^{\perp}$ almost surely. The claimed expression (27) follows immediately.
@@
Define the set $\mathcal{K}_{v}=\{\beta_{\star}\}^{\perp}\cup\{v\}^{\perp}$ for each unit vector $v\in\mathbb{S}^{d-1}$ and define the function $h\colon\mathbb{S}^{d-1}\rightarrow[0,\infty)$.
@@
Fatou’s lemma directly implies that $h$ is lower semicontinuous and therefore admits a minimizer $u_{\star}\in\operatornamewithlimits{argmin}_{u\in\mathbb{S}^{d-1}}h(u)$. Let $s\colon\mathbf{R}^{d}\rightarrow\mathbf{R}^{d}$ be a Borel measurable selection of the metric projection $\mathsf{P}_{\mathcal{K}_{u_{\star}}}$. Define the pushforward measure $\bar{\nu}=s_{\#}\mu$. Clearly $\bar{\nu}$ lies in $\mathcal{E}$ and hence
@@
If $(y_{1},y_{2})$ is a centered Gaussian vector with $\operatorname{\mathbb{E}}y_{1}^{2}=\sigma_{1}^{2}$, $\operatorname{\mathbb{E}}y_{2}^{2}=\sigma^{2}_{2}$, and $\operatorname{\mathbb{E}}y_{1}y_{2}=\rho\sigma_{1}\sigma_{2}$, then the equations hold:
$\operatorname{\mathbb{E}}|y_{1}y_{2}|$ $=\frac{2}{\pi}\left(\sqrt{1-\rho^{2}}+\rho\arcsin(\rho)\right)\sigma_{1}\sigma_{2},$
@@
The first equation is proved for example in [43, Corollary 3.1]. To see the second equation, standard results show that the conditional distribution $y_{1}\mid y_{2}$ is Gaussian $N(\rho(\sigma_{1}/\sigma_{2})y_{2},\sigma_{1}^{2}-\rho^{2}\sigma_{1}^{2})$. Thus, the second moment is $\operatorname{\mathbb{E}}[y_{1}^{2}\mid y_{2}]=(1-\rho^{2})\sigma_{1}^{2}+\frac{\rho^{2}\sigma_{1}^{2}}{\sigma_{2}^{2}}y_{2}^{2}$ and therefore iterating expectations gives $\operatorname{\mathbb{E}}y_{1}^{2}y_{2}^{2}=\operatorname{\mathbb{E}}[\operatorname{\mathbb{E}}[y_{1}^{2}\mid y_{2}]y_{2}^{2}]=(1+2\rho^{2})\sigma_{1}^{2}\sigma^{2}_{2},$ as claimed. ∎
@@
Consider the function $\psi\colon[-1,1]\rightarrow{\mathbf{R}}$ defined by $\psi(t)=\sqrt{1-t^{2}}+t\arcsin(t)$ and the function $\phi\colon[-1,1]\rightarrow{\mathbf{R}}$ given by $\phi(t)=\left(\frac{\pi}{2}-1\right)t^{2}+1$. Then, for any $t\in[-1,1]$ we have $\psi(t)\leq\phi(t).$

### v2

@@ Appendix C Proofs from Section 3
We first argue the inclusion $\supset$. Observe that for any measure $\mu\in\mathcal{P}_{2}^{\circ}$, the set of maximizers of (16) is the intersection of the unit sphere $\mathbb{S}^{d-1}$ and the top eigenspace of $\Sigma_{\mu}$. Thus, if $\lambda_{1}(\Sigma_{\mu})=\lambda_{2}(\Sigma_{\mu})$, then none of the maximizers are isolated, so they are unstable and therefore $\mu\in\mathcal{E}$.
To see the reverse inclusion $\subset$, fix a measure $\mu\in\mathcal{P}_{2}^{\circ}$ and suppose that the top two eigenvalues $\lambda_{1}$ and $\lambda_{2}$ of $\Sigma_{\mu}$ are distinct. Then, up to sign, the normalized top eigenvector $v_{\star}$ of $\Sigma_{\mu}$ is the unique maximizer of (16). It remains to verify that $v_{\star}$ is a tilt-stable maximizer of (16). To this end, define the Lagrangian function $\mathcal{L}(v,\lambda):=-\frac{1}{2}\operatorname{\mathbb{E}}_{\mu}\langle v,x\rangle^{2}+\frac{\lambda}{2}(\|v\|^{2}_{2}-1)$. Then the following equalities hold:
@@
In particular, the first equality shows that the optimal Lagrange multiplier $\lambda_{\star}$ is $\lambda_{1}$. Let $V\in\mathbf{R}^{d\times(d-1)}$ be a matrix with an orthonormal $\Sigma_{\mu}$-eigenbasis for $v_{\star}^{\perp}$ as its columns. Then an elementary computation yields
$\min_{y\in\mathbb{S}^{d-2}}\langle\nabla^{2}_{vv}\mathcal{L}(v_{\star},\lambda_{\star})(Vy),Vy\rangle=\min_{y\in\mathbb{S}^{d-2}}\sum_{i=2}^{d}(\lambda_{1}-\lambda_{i})y_{i-1}^{2}=\lambda_{1}-\lambda_{2}>0.$ (94)
Therefore $v_{\star}$ is a tilt-stable maximizer of (16) and hence $\mu\notin\mathcal{E}$. Moreover, the claimed equality ${\rm REG}(\mu)^{-1}=\lambda_{1}-\lambda_{2}$ follows directly from (94), thereby completing the proof.
@@ C.2 Proof of Theorem 3
This follows directly by applying Theorem A.2 with $\mathcal{G}=\mathcal{E}$ from Lemma 3.1 and $G={\{v\in\mathbf{R}^{d}_{+}\mid v^{(1)}=v^{(2)}\}}$, where $v^{(i)}$ denotes the $i$’th largest coordinate value of $v$.
@@
Henceforth, fix a measure $\mu\in\mathcal{P}_{2}^{\circ}$ and define the shorthand $\lambda:=\lambda(\Sigma_{\mu})$. Let’s dispense first with the simple direction $\supset$. To this end, suppose that equality $\lambda_{q}=\lambda_{q+1}$ holds. Then we may select two orthonormal sets $\mathcal{U}:=\{u_{1},\ldots,u_{q}\}$ and $\mathcal{U}^{\prime}:=\{u_{1},\ldots,u_{q-1},u_{q}^{\prime}\}$ with $\langle u_{q},u_{q}^{\prime}\rangle=0$ that are contained in the span of the eigenspaces corresponding to the top $q$ eigenvalues. We may further interpolate between these two sets via $\mathcal{U}_{t}:=\{u_{1},\ldots,tu_{q}+(1-t)u_{q}^{\prime}\}$ for $t\in[0,1]$. The orthogonal projections onto the span of each $\mathcal{U}_{t}$ furnish a continuous path of distinct maximizers of (19), which are therefore not tilt-stable. Thus $\mu\in\mathcal{E}$, as claimed.
@@
We now establish the reverse inclusion $\subset$. Suppose therefore that $\lambda_{q}$ and $\lambda_{q+1}$ are distinct. We begin by conveniently parameterizing the Grassmannian manifold ${\rm Gr}(q,d)$ as follows. Define the $d\times d$ matrix $A:=\begin{bmatrix}I_{q}&0\\ 0&0\end{bmatrix}$. Then, as in [6, Section 2.1], we may write ${\rm Gr}(q,d)$ as the orbit of $A$ under conjugation by orthogonal matrices:
${\rm Gr}(q,d)=\{UAU^{\top}\mid U\in O(d)\}.$
@@
Fix a skew-symmetric $d\times d$ matrix $V:=\begin{bmatrix}V_{1}&V_{2}\\ -V_{2}^{\top}&V_{4}\end{bmatrix}$, where $V_{2}\in\mathbf{R}^{q\times(d-q)}$, and define the curve $\gamma\colon\mathbf{R}\to{\rm Gr}(q,d)$ by
$\gamma(t):=\exp(-tV)\cdot A\cdot\exp(tV).$
Differentiating the curve $\gamma$ yields the expression
$\dot{\gamma}(0)=AV-VA=\begin{bmatrix}0&V_{2}\\ V_{2}^{\top}&0\end{bmatrix}\!$.
@@
Moreover, [6, Section 2.3] shows that varying $V$ among all skew-symmetric $d\times d$ matrices yields the entire tangent space at $A$:
$T_{{\rm Gr}(q,d)}(A)=\left\{\begin{bmatrix}0&B\\ B^{\top}&0\end{bmatrix}~\Bigg|~B\in\mathbf{R}^{q\times(d-q)}\right\}\!.$
@@
Without loss of generality, we may assume that $\Sigma_{\mu}$ is diagonal with $\Sigma_{\mu}={\rm Diag}(\lambda)$. Then $A$ is clearly the unique maximizer of the problem (19). Consider now the second-order expansion
$(f\circ\gamma)(t)=\langle\gamma(t),\Sigma_{\mu}\rangle=f(A)+t\underbrace{\langle AV-VA,\Sigma_{\mu}\rangle}_{=0}+\tfrac{1}{2}t^{2}\langle AV^{2}+V^{2}A-2VAV,\Sigma_{\mu}\rangle+O(t^{3}).$
In particular,
$\nabla^{2}_{{\rm Gr}(q,d)}f(A)[\dot{\gamma}(0),\dot{\gamma}(0)]=(f\circ\gamma)^{\prime\prime}(0)=\langle AV^{2}+V^{2}A-2VAV,\Sigma_{\mu}\rangle.$
Taking into account the definition of $A$, a quick computation shows
$AV^{2}+V^{2}A-2VAV$ $=\begin{bmatrix}2(V_{1}^{2}-V_{2}V_{2}^{\top})&V_{1}V_{2}+V_{2}V_{4}\\ (V_{1}V_{2}+V_{2}V_{4})^{\top}&0\end{bmatrix}-2\begin{bmatrix}V_{1}^{2}&V_{1}V_{2}\\ (V_{1}V_{2})^{\top}&-V_{2}^{\top}V_{2}\end{bmatrix}$
$=\begin{bmatrix}-2V_{2}V_{2}^{\top}&V_{2}V_{4}-V_{1}V_{2}\\ (V_{2}V_{4}-V_{1}V_{2})^{\top}&2V_{2}^{\top}V_{2}\end{bmatrix}\!$.
@@
Then taking the trace product with the diagonal matrix $\Sigma_{\mu}={\rm Diag}(\lambda)$ yields
$(f\circ\gamma)^{\prime\prime}(0)$ $=-2\langle V_{2}V_{2}^{\top},{\rm Diag}(\lambda_{1:q})\rangle+2\langle V_{2}^{\top}V_{2},{\rm Diag}(\lambda_{q+1:d})\rangle$ (95)
$\leq-2\langle V_{2}V_{2}^{\top},\lambda_{q}I_{q}\rangle+2\langle V_{2}^{\top}V_{2},\lambda_{q+1}I_{d-q}\rangle$ (96)
$=-(\lambda_{q}-\lambda_{q+1})\cdot 2\|V_{2}\|^{2}_{F}$ (97)
$=-(\lambda_{q}-\lambda_{q+1})\|\dot{\gamma}(0)\|^{2}_{F}.$ (98)
@@
Consequently, the covariant Hessian $\nabla^{2}_{{\rm Gr}(q,d)}f(A)$ is negative definite on the tangent space $T_{{\rm Gr}(q,d)}(A)$. Therefore $A$ is a tilt-stable maximizer of the problem (19) and hence $\mu\notin\mathcal{E}$. Moreover, if we take $V_{2}$ to be any scalar multiple of the $q\times(d-q)$ matrix with $(q,1)$’st entry $1$ and all other entries 0, then (96) holds with equality. Hence ${\rm REG}(\mu)^{-1}=\lambda_{q}-\lambda_{q+1}$, as claimed.
@@ C.4 Proof of Theorem 3
This follows directly by applying Theorem A.2 with $\mathcal{G}=\mathcal{E}$ from Lemma 3.3 and $G={\{v\in\mathbf{R}^{d}_{+}\mid v^{(q)}=v^{(q+1)}\}}$, where $v^{(i)}$ denotes the $i$’th largest coordinate value of $v$.
@@ Appendix D Proofs from Section 4
First, observe
$\nabla f(\beta)={\operatorname{\mathbb{E}}}\big[(h^{\prime}(\langle x,\beta\rangle)-y)x\big]$ $={\operatorname{\mathbb{E}}}\big[{\operatorname{\mathbb{E}}}\big[h^{\prime}(\langle x,\beta\rangle)-y\!\mid\!x\big]x\big]$
$={\operatorname{\mathbb{E}}}\big[(h^{\prime}(\langle x,\beta\rangle)-h^{\prime}(\langle x,{\beta}_{\star}\rangle))x\big].$
Therefore, the equality $\nabla f({\beta}_{\star})=0$ holds and hence $\beta$$_{\star}$ is a critical point for the problem with optimal Lagrange multipliers $\lambda_{\star}=0$. Differentiating again yields the expression for the Hessian
$H:=\nabla^{2}f({\beta}_{\star})={\operatorname{\mathbb{E}}}\big[h^{\prime\prime}(\langle x,\beta_{\star}\rangle)xx^{\top}\big].$ (99)
Note that $H$ is positive semidefinite since $h^{\prime\prime}>0$. Consequently, the set of ill-conditioned distributions $\mathcal{E}$ corresponds to those distributions $\mu\in\mathcal{Q}$ of $x$ for which $\ker(H)$ nontrivially intersects $\mathcal{T}$. Clearly, a vector $v$ lies in $\ker(H)$ if and only if $\langle Hv,v\rangle=0$, or equivalently ${\operatorname{\mathbb{E}}}\big[h^{\prime\prime}(\langle x,{\beta}_{\star}\rangle)\langle x,v\rangle^{2}\big]=0$. Taking into account the assumption $h^{\prime\prime}>0$, this occurs precisely when $v$ lies in the nullspace of $\Sigma_{\mu}$. Thus $\mathcal{E}$ is comprised of all measures $\mu\in\mathcal{Q}$ satisfying $\mathcal{T}\cap\ker(\Sigma_{\mu})\neq\{0\}.$ Finally, it follows directly from (99) that if for some $a,b>0$ the inequalities $a\leq h^{\prime\prime}(\langle x,\beta_{\star}\rangle)\leq b$ hold for $\mu$-almost every $x$, then $\lambda_{\min}(H|_{{\mathcal{T}}})\in\lambda_{\min}(\Sigma_{\mu}|_{{\mathcal{T}}})\cdot[a,b]$, thereby completing the proof.
@@ D.2 Proof of Theorem 4
We begin by verifying the claim for the formulation (28). To this end, a quick computation shows $\nabla f(\beta_{\star})=0$ and $\nabla^{2}f(\beta_{\star})=\operatorname{\mathbb{E}}_{\mu}\langle x,{\beta}_{\star}\rangle^{2}xx^{\top}=\hat{\Sigma}_{\mu},$ where $\mu=\mathsf{D}_{x}$. Therefore we deduce
@@
In particular, $\nabla^{2}f(\beta_{\star})$ is singular if and only if the support of $\mu$ is contained in ${\beta}_{\star}^{\perp}\cup v^{\perp}$ for some $v\in\mathbb{S}^{d-1}$. If $\mu\in\mathcal{P}_{4}\setminus\mathcal{E}$, then clearly (29) holds with $(c_{\textup{{lb}}},c_{\textup{{ub}}})=(1,1)$.
Next, we verify the claim for the formulation (27). To this end, let $\mathcal{M}$ denote the manifold of symmetric PSD rank-one matrices:
$\mathcal{M}=\{M\in\bm{S}^{d}_{+}\mid\operatorname{rank}(M)=1\}.$
A quick computation yields
$\nabla f(M)={\operatorname{\mathbb{E}}_{\mu}}\big\langle M-M_{\star},xx^{\top}\big\rangle xx^{\top}\qquad\textrm{and}\qquad\nabla^{2}f(M_{\star})[\Delta,\Delta]={\operatorname{\mathbb{E}}_{\mu}}\big\langle\Delta,xx^{\top}\big\rangle^{2}.$
In particular, equality $\nabla f(M_{\star})=0$ holds and therefore the optimal Lagrange multipliers $\lambda_{\star}$ are zero. Hence the Hessian of the Lagrangian at $(M_{\star},\lambda_{\star})$ coincides with $\nabla^{2}f(M_{\star})$. Classically, if we form the factorization $M_{\star}={\beta}_{\star}{\beta}_{\star}^{\top}$, then the tangent space $\mathcal{T}$} to $\mathcal{M}$ at $M_{\star}$ can be written as
$\mathcal{T}=\{{\beta}_{\star}v^{\top}+v{\beta}_{\star}^{\top}\mid v\in\mathbf{R}^{d}\}.$
@@
$\nabla^{2}f(M_{\star})[\Delta,\Delta]={\operatorname{\mathbb{E}}_{\mu}}\big\langle\Delta,xx^{\top}\big\rangle^{2}=4\operatorname{\mathbb{E}}_{\mu}\langle{\beta}_{\star},x\rangle^{2}\langle v,x\rangle^{2}=4\langle\hat{\Sigma}_{\mu}v,v\rangle.$
Note $\|\Delta\|_{F}^{2}=2\langle{\beta}_{\star},v\rangle^{2}+2\|{\beta}_{\star}\|^{2}_{2}\|v\|^{2}_{2}$ and therefore $\|\Delta\|_{F}^{2}/\|{\beta}_{\star}\|^{2}_{2}\|v\|^{2}_{2}\in[2,4]$ whenever $v$ is nonzero. Hence $\Delta$ is nonzero whenever $v$ is nonzero. We therefore again deduce $\nabla^{2}f(M_{\star})[\Delta,\Delta]=0$ for some nonzero $\Delta\in{\mathcal{T}}$ if and only if the support of $\mu$ is contained in ${\beta}_{\star}^{\perp}\cup v^{\perp}$ for some $v\in\mathbb{S}^{d-1}$. Finally, if $\mu\in\mathcal{P}_{4}\setminus\mathcal{E}$, then (29) follows immediately with $(c_{\textup{{lb}}},c_{\textup{{ub}}})=\big(\tfrac{1}{2}\|{\beta}_{\star}\|^{2}_{2},\|{\beta}_{\star}\|^{2}_{2}\big)$.
@@
Define the set $\mathcal{K}_{v}:=\beta_{\star}^{\perp}\cup v^{\perp}$ for each unit vector $v\in\mathbb{S}^{d-1}$ and let $h\colon\mathbb{S}^{d-1}\rightarrow[0,\infty)$ be given by
@@
Fatou’s lemma directly implies that $h$ is lower semicontinuous and therefore admits a minimizer $u_{\star}\in\operatornamewithlimits{argmin}_{u\in\mathbb{S}^{d-1}}h(u)$. Let $s\colon\mathbf{R}^{d}\rightarrow\mathcal{K}_{u_{\star}}$ be a Borel measurable selection of the metric projection $\mathsf{P}_{\mathcal{K}_{u_{\star}}}$. Define the pushforward measure $\bar{\nu}:=s_{\#}\mu$. Clearly $\bar{\nu}$ lies in $\mathcal{E}$ and hence
@@
If $(y_{1},y_{2})$ is a centered Gaussian vector with $\operatorname{\mathbb{E}}y_{1}^{2}=\sigma_{1}^{2}$, $\operatorname{\mathbb{E}}y_{2}^{2}=\sigma^{2}_{2}$, and $\operatorname{\mathbb{E}}y_{1}y_{2}=\rho\sigma_{1}\sigma_{2}$ for some $\sigma_{1},\sigma_{2}>0$ and $\rho\in[-1,1]$, then the following equalities hold:
$\operatorname{\mathbb{E}}|y_{1}y_{2}|$ $=\frac{2}{\pi}\bigg(\sqrt{1-\rho^{2}}+\rho\arcsin(\rho)\bigg)\sigma_{1}\sigma_{2},$
@@
The first equation is proved for example in [45, Corollary 3.1]. To see the second equation, standard results show $y_{1}\!\mid\!y_{2}\sim\mathsf{N}\big(\rho(\sigma_{1}/\sigma_{2})y_{2},(1-\rho^{2})\sigma_{1}^{2}\big)$. Hence $\operatorname{\mathbb{E}}[y_{1}^{2}\!\mid\!y_{2}]=(1-\rho^{2})\sigma_{1}^{2}+\frac{\rho^{2}\sigma_{1}^{2}}{\sigma_{2}^{2}}y_{2}^{2}$, and since $\operatorname{\mathbb{E}}y_{2}^{4}=3\sigma_{2}^{4}$, the tower rule now implies $\operatorname{\mathbb{E}}(y_{1}y_{2})^{2}={\operatorname{\mathbb{E}}}\big[{\operatorname{\mathbb{E}}}[y_{1}^{2}\!\mid\!y_{2}]y_{2}^{2}\big]=(1+2\rho^{2})\sigma_{1}^{2}\sigma^{2}_{2}$. ∎
@@
Consider the function $\psi\colon[-1,1]\rightarrow{\mathbf{R}}$ given by $\psi(t)=\sqrt{1-t^{2}}+t\arcsin(t)$ and the function $\phi\colon[-1,1]\rightarrow{\mathbf{R}}$ given by $\phi(t)=\big(\frac{\pi}{2}-1\big)t^{2}+1$. Then $\psi(t)\leq\phi(t)$ for all $t\in[-1,1]$.